OpenStudy (kamille):

i need to solve \[\log _{0,7}\log _{2}\frac{ 3+5x }{x ^{2}+3}<\log _{0,7}1\]

OpenStudy (kamille):

@Mertsj

OpenStudy (kamille):

@kausarsalley do you have any tips?

OpenStudy (anonymous):

no...but i want to try it out...

OpenStudy (kamille):

@ash2326

OpenStudy (anonymous):

@Mertsj

OpenStudy (kamille):

Cant it be written as \[\[\log _{0,7}\log _{2}\frac{ 3+5x }{x ^{2}+3}<\log _{0,7}\log _{2}2\]\]

OpenStudy (kamille):

?

OpenStudy (kamille):

@hartnn

OpenStudy (kamille):

@AravindG

OpenStudy (kamille):

@saifoo.khan

OpenStudy (mertsj):

Well, first of all, the log of 1 is 0 so the right side is 0

OpenStudy (kamille):

oh,okay

OpenStudy (mertsj):

Is that base 0.7?

OpenStudy (kamille):

yes

OpenStudy (kamille):

it is a loghorhytm in logorhytm

OpenStudy (kamille):

actually, I have made to this logorhytm. the problem is: \[5^{\log_{0,7}\log_{2}\frac{3+5x}{x^{2}+3}}<1\]

OpenStudy (kamille):

maybe I have done something wrong!

OpenStudy (kamille):

@Mertsj any ideas?

OpenStudy (kamille):

well, let me show how I done, maybe you will find my mistake,okay?

OpenStudy (kamille):